get the value of x and y
x+3y=9
2x-y=4

Answer by checkley77(12844)

x+3y=9
2x-y=4 multiply by 3 & add.
6x-3y=12
x+3y=9
---------------
7x=21
x=21/7
x=3 ans.
3+3y=9
3y=9-3
3y=6
y=6/3
y=2 ans.
Proof:
2*3-2=4
6-2=4
4=4